OpenStudy (anonymous):

Use the half angle formula to find exact value of cos(3pi/8)

myininaya (myininaya):

so it it says to use half angle formula... We know \[\cos^2(\frac{x}{2})=\frac{1}{2}(1+\cos(x))\]

myininaya (myininaya):

\[\frac{3 \pi}{8}=\frac{x}{2}\] What will you be pluggin in for x?